OpenStudy (kj4uts):

For a sample size of n=26 and a population parameter of p=0.4, a normal curve can be used to approximate the binomial probability histogram. True or false?

OpenStudy (perl):

you have to check that n*p >= 5 and n*(1-p)>= 5

OpenStudy (kj4uts):

@perl np = 26 * 0.4 = 10.4 > 5 would this be false

OpenStudy (amistre64):

check 26(.4) and 26(.6) if they are both bigger than 5 you should be fine

OpenStudy (amistre64):

2.6 = 10% they are fine

OpenStudy (kj4uts):

@amistre64 so it is false?